# Raw Data Matters: Enhancing Prompt Tuning by Internal Augmentation on Vision-Language Models

### v1

@@ A.1 Experimental Setup
As a distillation-based approach, the data sampling strategy for the fine-tuning and inference process of AugPT remains consistent with the backbone model PromptKD (Li et al. 2024a).
@@
[…]
@@ A.2 More Details of Distillation-based Prompt Tuning Backbone
AugPT introduces PromptSRC (Khattak et al. 2023b) pre-tuned by PromptKD using ViT-L/14 encoders as a teacher model. As a supplement to Preliminaries section (Sec. 3.1) in main text, we provide more detailed description of the teacher model and the fine-tuning process associated with the standard paradigm of Distillation-based Prompt Tuning approaches to which AugPT belongs.
@@
PromptSRC adopts the IVLP (Rasheed et al. 2023) (Independent Visual-Language Prompting) framework by separately optimizing visual and textual prompts during fine-tuning. To alleviate the performance degradation caused by the Base-to-New Trade-off (BNT) problem (Zhou et al. 2022a; Zhang et al. 2024a; Li et al. 2025a), it introduces a series of structurally aligned regularization losses. Beyond the conventional cross-entropy loss $\mathcal{L}_{\text{CE}}$ commonly imported in mainstream prompt tuning methods, PromptSRC enforces additional consistency constraints across modalities, which are defined as $\mathcal{L}_{\text{SCL-image}}$ and $\mathcal{L}_{\text{SCL-text}}$, encouraging better alignment between prompts and their respective modality-specific features, while $\mathcal{L}_{\text{SCL-logits}}$ further regularizes the logits after cross-modal interaction. These jointly contribute to improved generalization across both base and new tasks. The framework of PromptSRC is illustrated in Fig. 4. The optimizing goal of fine-tuning can be summarized as follows: by continuously optimizing the weighted sum of the cross-entropy loss $\mathcal{L}_{\text{CE}}$ and the consistency loss $\mathcal{L}_{\text{SCL}}$ , the output logits obtained from the learnable prompts as input queries are progressively aligned with the one-hot distribution of the ground-truth label.
@@
[…]
@@
[…]
The fine-tuning method for the student model is described in Preliminaries section (Sec. 3.1). The shape of 9 student image prompts are [4,512][4,512][ 4 , 512 ]. AugPT further introduces an optimized variant of this procedure in Method section (Sec. 3.4) in main text, referred to as Optimized Prompt Distillation (OPD).
@@ A.3 More Details of Adaptive Self-supervised Augmentation
AugPT follows the setup of RandAugment (Cubuk et al. 2020) by independently sampling $S$ augmentation operations from a pre-defined set of mainstream image transformation policies $\mathcal{Y}=\{\gamma_{1},\dots,\gamma_{Q}\}$ for constructing a dynamic policy group $\mathcal{G}$ on a single augmentation, then performs $N$ self-supervised augmentation on the raw unlabeled image $I^{\prime}$ to obtain the augmented set $\mathcal{D}(I^{\prime})$ with $N+1$ views. The complete list of $Q=16$ candidate augmentation operations along with their corresponding magnitude ranges $(A_{\text{min}},A_{\text{max}})$ is provided in Tab. 7.
@@
For the fixed amplitude hyperparameter $A$ pre-searched by RandAugment (on CIFAR10 (Krizhevsky, Hinton et al. 2009)), the transformation strength for each image operation $\hat{A}$ is determined by the following conversion:
@@ A.4 Pseudo-code of Algorithm
zero-shot CLIP (Radford et al. 2021) 69.34 74.22 71.70
few-shot CoCoOp (Zhou et al. 2022a) 80.47 71.69 75.83
KgCoOp (Yao, Zhang, and Xu 2023) 80.73 73.60 77.00
MaPLe (Khattak et al. 2023a) 82.28 75.14 78.55
TCP (Yao, Zhang, and Xu 2024) 84.13 75.36 79.51
MMRL (Guo and Gu 2025) 85.68 77.16 81.20
@@
MAO (Li et al. 2025b) 84.53 75.38 79.69
CasPL (Wu et al. 2024) 86.11 79.54 82.69
unlabeled images KDPL (Mistretta et al. 2024) 77.11 71.61 74.26
@@ B.1 Compare with More Baselines
To further compare the performance of AugPT against a wider range of prompt tuning models, we introduce additional baseline models in Tab. 8. These models include the original zero-shot CLIP (Radford et al. 2021), as well as CoCoOp (Zhou et al. 2022a), KgCoOp (Yao, Zhang, and Xu 2023), MaPLe (Khattak et al. 2023a), TCP (Yao, Zhang, and Xu 2024), and MMRL (Guo and Gu 2025), which are all fine-tuned with few-shot image-text pairs. Additionally, we incorporate MAO (Li et al. 2025b) and CasPL (Wu et al. 2024), which utilize a combination of few-shot image-text pairs and unlabeled images. Another distillation-based prompt tuning approach KDPL (Mistretta et al. 2024) is also contained in the comparison.
It can be observed that AugPT requires only 16-shot unlabeled images to achieve performance comparable to advanced prompt tuning models (Guo and Gu 2025) fine-tuned with the same shot of image-text pairs, indicating a lower data requirement. Similarly, when fine-tuned on full dataset, AugPT outperforms models (Wu et al. 2024) that use both few-shot image-text pairs and full unlabeled images.
@@
Table 10: Detailed results of ablation under different augmented images $N$ on full base-to-new tasks.
@@ B.2 Error Bar Analysis
(2) $N$ Random Resized Crop 83.72 77.44 44.94 68.70
@@ B.3 Effectiveness of Adaptive Self-supervised Augmentation
To further evaluate the performance trend of the ASA module (Method section in main text) in AugPT under different numbers of augmented images $N$, we provide the detailed numerical results corresponding to Fig. 5(a) in Tab. 10.
@@ B.4 Effectiveness of Consensus-based Filtering Gate
Specifically, we replace the pre-tuned teacher model used for online inference in the Consensus Test (i.e., PromptSRC (Khattak et al. 2023b) with ViT-L/14 encoders) with the zero-shot foundation CLIP model without fine-tuning. Since the foundation CLIP has no knowledge of the in-domain data distribution, its ability to infer consensus over image features in base tasks is expected to degrade. We experiment with both ViT-L/14 and ViT-B/16 versions of zero-shot CLIP as the Filtering Gate, and execute Top-1 Consensus Test for fair comparison.
@@
[…]
@@ B.5 Adaptability of AugPT
To further evaluate the adaptability of AugPT, we include a comparison with the UPL (Huang, Chu, and Wei 2022) framework, which shares a similar design philosophy. UPL is a prompt tuning method that also relies solely on unlabeled images, using the frozen pretrained CLIP model as supervision and guiding CoOp-based prompt tuning through the query results of images. For a fair comparison, when integrating AugPT into this setup, we similarly replace the teacher model and Filtering Gate with the foundation CLIP model. In addition, we follow the original UPL implementation, where model performance is measured by fine-tuning accuracy using the full label set as base classes. To prevent data leakage, test images are excluded from tuning process.
@@ B.6 Computational Cost
In Tab. 15, we evaluate the computational overhead of AugPT compared to the PromptKD backbone during base-class fine-tuning. Using the StanfordCars (Krause et al. 2013) dataset as an example, we report metrics including the number of learnable parameters, GPU memory usage, image processing speed measured by Frames Per Second (FPS) during fine-tuning, and HM performance for comparison.
We observe that AugPT possesses moderate increases in GPU memory usage and processing time, which can be attributed to the additional computational cost introduced by parallel-handled augmented views. Nonetheless, since AugPT does not attach more learnable parameters and avoids the overhead of collecting external knowledge, the overall increase in fine-tuning overhead remains relatively low. Moreover, since AugPT is not influenced by query-time bottleneck which is unavoidable in LLM-based prompt tuning methods, its fine-tuning FPS (36.1) has a significant advantage over the typical LLM-guided models (normally $<$10). Overall, compared to other data-driven prompt tuning approaches, the fine-tuning cost of AugPT is acceptable.
@@
For inference-stage computational cost, we compare FPS as the evaluation metric in Fig. 6, benchmarking AugPT against advanced data-driven prompt tuning models including HPT (Wang et al. 2024), LLaMP (Zheng et al. 2024), and PromptKD (Li et al. 2024a). It is evident that AugPT achieves faster inference speeds than LLM-guided models and basically remains consist with its backbone, PromptKD, while realizing higher HM performance. This demonstrates the parameter efficiency of AugPT during inference.
@@ Appendix D Limitations and Future Work
Although AugPT surpasses current SOTA method using only internal augmentation without introducing external knowledge or additional learnable parameters, we believe there is still room for improvement. Firstly, as mentioned in Sec. B.5, the design of AugPT is tightly coupled with distillation-based prompt tuning, making it less adaptable to other prompt tuning paradigms. Secondly, following most existing studies (Zhou et al. 2022b; Khattak et al. 2023b; Li et al. 2025a; Guo and Gu 2025) in this area, AugPT focuses on improving performance for recognition-related tasks, while its applicability to other format of tasks (e.g., semantic segmentation) remains unexplored. Finally, the parallel processing of augmented views introduces a moderate increase in computational cost. In future work, we plan to extend AugPT to support a broader range of backbone models and downstream tasks, and consider improving its computational efficiency.

### v2

@@ A.1 Experimental Setup
As a distillation-based approach, the data sampling strategy for the fine-tuning and inference process of AugPT remains consistent with the backbone model PromptKD (li2024promptkd).
@@
Following the setup of the PromptKD backbone, as the initialization setup, we set the the length of learnable text or image prompts in the teacher and student model to 4, as well as the depth to 9 (insert prompts between the first 9 Transformer blocks in ViT encoders). The learnable image prompts of teacher and student model are randomly initialized adhering to a zero-mean Gaussian distribution $X\sim\mathcal{N}(0,\ 0.02^{2})$. The learnable feature projection layer $\text{Proj}(\cdot)$ in student model is initialized randomly with 2 MLP layers (ablation study of layers is in Sec. B.4). The fine-tuning details of the applied teacher model are introduced in Sec. A.2.
@@ A.2 More Details of Distillation-based Prompt Tuning Backbone
AugPT introduces PromptSRC (khattak2023promptsrc) pre-tuned by PromptKD using ViT-L/14 encoders as a teacher model. As a supplement to Preliminaries section (Sec. 3.1) in main text, we provide more detailed description of the teacher model and the fine-tuning process associated with the standard paradigm of Distillation-based Prompt Tuning approaches to which AugPT belongs.
@@
PromptSRC adopts the IVLP (rasheed2023ivlp) (Independent Visual-Language Prompting) framework by separately optimizing visual and textual prompts during fine-tuning. To alleviate the performance degradation caused by the Base-to-New Trade-off (BNT) problem (zhou2022cocoop; zhang2024dept; li2025dpc), it introduces a series of structurally aligned regularization losses. Beyond the conventional cross-entropy loss $\mathcal{L}_{\text{CE}}$ commonly imported in mainstream prompt tuning methods, PromptSRC enforces additional consistency constraints across modalities, which are defined as $\mathcal{L}_{\text{SCL-image}}$ and $\mathcal{L}_{\text{SCL-text}}$, encouraging better alignment between prompts and their respective modality-specific features, while $\mathcal{L}_{\text{SCL-logits}}$ further regularizes the logits after cross-modal interaction. These jointly contribute to improved generalization across both base and new tasks. The framework of PromptSRC is illustrated in Fig. 4. The optimizing goal of fine-tuning can be summarized as follows: by continuously optimizing the weighted sum of the cross-entropy loss $\mathcal{L}_{\text{CE}}$ and the consistency loss $\mathcal{L}_{\text{SCL}}$ , the output logits obtained from the learnable prompts as input queries are progressively aligned with the one-hot distribution of the ground-truth label.
@@
In the fine-tuning stage, serving as the teacher model in distillation-based prompt tuning approaches, PromptSRC (khattak2023promptsrc) adopts a pre-trained CLIP model with larger ViT-L/14 encoders as the foundation model. At initialization, PromptSRC uses the prompt template “A photo of a [CLS]” to initialize the learnable text prompts and samples the learnable image prompts from a zero-mean Gaussian distribution $X\sim\mathcal{N}(0,\ 0.02^{2})$. Both text and image prompts have a depth of 9 and a shape of $[4,768]$.
@@
where $\text{sim}[\cdot,\cdot]$ denotes cosine similarity and $\tau$ is the temperature coefficient. It is worth noting that, following the PromptKD (li2024promptkd) setting, the candidate set $\boldsymbol{C}$ includes all class names, containing both base and new classes. Since only unlabeled images are introduced during fine-tuning as raw data, the relationship between class labels and images is entirely unknown in the fine-tuning process, ensuring that no data leakage occurs.
The fine-tuning method for the student model is described in Preliminaries section (Sec. 3.1). The shape of 9 student image prompts are $[4,512]$. AugPT further introduces an optimized variant of this procedure in Method section (Sec. 3.4) in main text, referred to as Optimized Prompt Distillation (OPD).
@@ A.3 More Details of Adaptive Self-supervised Augmentation
AugPT follows the setup of RandAugment (cubuk2020randaugment) by independently sampling $S$ augmentation operations from a pre-defined set of mainstream image transformation policies $\mathcal{Y}=\{\gamma_{1},\dots,\gamma_{Q}\}$ for constructing a dynamic policy group $\mathcal{G}$ on a single augmentation, then performs $N$ self-supervised augmentation on the raw unlabeled image $I^{\prime}$ to obtain the augmented set $\mathcal{D}(I^{\prime})$ with $N+1$ views. The complete list of $Q=16$ candidate augmentation operations along with their corresponding magnitude ranges $(A_{\text{min}},A_{\text{max}})$ is provided in Tab. 7.
@@
For the fixed amplitude hyperparameter $A$ pre-searched by RandAugment (on CIFAR10 (krizhevsky2009cifar10)), the transformation strength for each image operation $\hat{A}$ is determined by the following conversion:
@@ A.4 Pseudo-code of Algorithm
zero-shot CLIP (radford2021clip) 69.34 74.22 71.70
few-shot CoCoOp (zhou2022cocoop) 80.47 71.69 75.83
KgCoOp (yao2023kgcoop) 80.73 73.60 77.00
MaPLe (khattak2023maple) 82.28 75.14 78.55
TCP (yao2024tcp) 84.13 75.36 79.51
MMRL (guo2025mmrl) 85.68 77.16 81.20
@@
MAO (li2025mao) 84.53 75.38 79.69
CasPL (wu2024caspl) 86.11 79.54 82.69
unlabeled images KDPL (mistretta2024kdpl) 77.11 71.61 74.26
@@ B.1 Compare with More Baselines
To further compare the performance of AugPT against a wider range of prompt tuning models, we introduce additional baseline models in Tab. 8. These models include the original zero-shot CLIP (radford2021clip), as well as CoCoOp (zhou2022cocoop), KgCoOp (yao2023kgcoop), MaPLe (khattak2023maple), TCP (yao2024tcp), and MMRL (guo2025mmrl), which are all fine-tuned with few-shot image-text pairs. Additionally, we incorporate MAO (li2025mao) and CasPL (wu2024caspl), which utilize a combination of few-shot image-text pairs and unlabeled images. Another distillation-based prompt tuning approach KDPL (mistretta2024kdpl) is also contained in the comparison.
It can be observed that AugPT requires only 16-shot unlabeled images to achieve performance comparable to advanced prompt tuning models (guo2025mmrl) fine-tuned with the same shot of image-text pairs, indicating a lower data requirement. Similarly, when fine-tuned on full dataset, AugPT outperforms models (wu2024caspl) that use both few-shot image-text pairs and full unlabeled images.
@@
Table 10: Detailed results of ablation under different augmented images NN on full base-to-new tasks.
@@ B.2 Error Bar Analysis
(2) NN Random Resized Crop 83.72 77.44 44.94 68.70
@@ B.3 Effectiveness of Adaptive Self-supervised Augmentation
To further evaluate the performance trend of the ASA module (Method section in main text) in AugPT under different numbers of augmented images NN, we provide the detailed numerical results corresponding to Fig. 5(a) in Tab. 10.
@@ B.4 Effectiveness of Consensus-based Filtering Gate
Specifically, we replace the pre-tuned teacher model used for online inference in the Consensus Test (i.e., PromptSRC (khattak2023promptsrc) with ViT-L/14 encoders) with the zero-shot foundation CLIP model without fine-tuning. Since the foundation CLIP has no knowledge of the in-domain data distribution, its ability to infer consensus over image features in base tasks is expected to degrade. We experiment with both ViT-L/14 and ViT-B/16 versions of zero-shot CLIP as the Filtering Gate, and execute Top-1 Consensus Test for fair comparison.
@@
In Optimized Prompt Distillation, AugPT follows the PromptKD (li2024promptkd) backbone by introducing an MLP-based projection layer $\text{Proj}(\cdot)$ to align student image features with teacher text features. To further verify the potential impact of the number of layers in this projection module, we conduct an ablation study in Tab. 13 by varying the default 2-layer MLP to 1 or 3 layers.
@@ B.5 Adaptability of AugPT
To further evaluate the adaptability of AugPT, we include a comparison with the UPL (huang2022upl) framework, which shares a similar design philosophy. UPL is a prompt tuning method that also relies solely on unlabeled images, using the frozen pretrained CLIP model as supervision and guiding CoOp-based prompt tuning through the query results of images. For a fair comparison, when integrating AugPT into this setup, we similarly replace the teacher model and Filtering Gate with the foundation CLIP model. In addition, we follow the original UPL implementation, where model performance is measured by fine-tuning accuracy using the full label set as base classes. To prevent data leakage, test images are excluded from tuning process.
@@ B.6 Computational Cost
In Tab. 15, we evaluate the computational overhead of AugPT compared to the PromptKD backbone during base-class fine-tuning. Using the StanfordCars (krause2013cars) dataset as an example, we report metrics including the number of learnable parameters, GPU memory usage, image processing speed measured by Frames Per Second (FPS) during fine-tuning, and HM performance for comparison.
We observe that AugPT possesses moderate increases in GPU memory usage and processing time, which can be attributed to the additional computational cost introduced by parallel-handled augmented views. Nonetheless, since AugPT does not attach more learnable parameters and avoids the overhead of collecting external knowledge, the overall increase in fine-tuning overhead remains relatively low. Moreover, since AugPT is not influenced by query-time bottleneck which is unavoidable in LLM-based prompt tuning methods, its fine-tuning FPS (36.1) has a significant advantage over the typical LLM-guided models (normally <<10). Overall, compared to other data-driven prompt tuning approaches, the fine-tuning cost of AugPT is acceptable.
@@
For inference-stage computational cost, we compare FPS as the evaluation metric in Fig. 6, benchmarking AugPT against advanced data-driven prompt tuning models including HPT (wang2024hpt), LLaMP (zheng2024llamp), and PromptKD (li2024promptkd). It is evident that AugPT achieves faster inference speeds than LLM-guided models and basically remains consist with its backbone, PromptKD, while realizing higher HM performance. This demonstrates the parameter efficiency of AugPT during inference.
@@ Appendix D Limitations and Future Work
Although AugPT surpasses current SOTA method using only internal augmentation without introducing external knowledge or additional learnable parameters, we believe there is still room for improvement. Firstly, as mentioned in Sec. B.5, the design of AugPT is tightly coupled with distillation-based prompt tuning, making it less adaptable to other prompt tuning paradigms. Secondly, following most existing studies (zhou2022coop; khattak2023promptsrc; li2025dpc; guo2025mmrl) in this area, AugPT focuses on improving performance for recognition-related tasks, while its applicability to other format of tasks (e.g., semantic segmentation) remains unexplored. Finally, the parallel processing of augmented views introduces a moderate increase in computational cost. In future work, we plan to extend AugPT to support a broader range of backbone models and downstream tasks, and consider improving its computational efficiency.